OpenStudy (anonymous):

Solve for x: [2(a-x)/3cube root of x]-x^2/3=0

OpenStudy (anonymous):

\[\frac{ 2(a-x) }{3\sqrt[3]{x} }-x ^{2/3}=0\] So this is basically what equation look like, and solve for x

OpenStudy (lgbasallote):

try multiplying EVERYTHING by \(\large \sqrt[3]x\)

OpenStudy (anonymous):

Sorry, I'm kinda stuck there. I'm try think, but what do I get by multiplying everything by \[\sqrt[3]{x}\]

OpenStudy (lgbasallote):

hmm sorry...i got confused now too

OpenStudy (lgbasallote):

i know...LCD \[\huge \implies \frac{2(a-x)}{3\sqrt[3] x} - \frac{3x}{\sqrt[3]x} = 0\]

OpenStudy (lgbasallote):

the denominator of the fraction should have a3

OpenStudy (lgbasallote):

\[\huge \frac{2(a-x)}{3\sqrt[3]x} - \frac{3x}{3\sqrt[3]x} = 0\]

OpenStudy (anonymous):

yeah, I got to that first part.

OpenStudy (lgbasallote):

now combine them \[\huge \implies \frac{2(a-x) -3x}{3\sqrt[3]x} = 0\] now...cross multiply..what do you get?

OpenStudy (anonymous):

2a-5x=0

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so solve for x

OpenStudy (anonymous):

2a/5

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so x = 2a/5

OpenStudy (anonymous):

ok, awesome...thanks

OpenStudy (lgbasallote):

welcome